[Prev][Next][Index][Thread]

(meteorobs) Fwd: BULLETIN: First Images from the SOHO Spacecraft - Status Report




-- BEGIN included message

                        SOHO SPACECRAFT STATUS REPORT

                          First Images Look GREAT !!

                               14 October 1998

     The Solar and Heliospheric Observatory (SOHO) spacecraft is performing
above expectations since control of the spacecraft was recently regained.
Recommissioning activities for the numerous scientific instruments are
continuing and will persist through the remainder of October.

     Over the weekend and today, the MDI/SOI (Michelson Doppler Imager/Solar
Oscillations Investigation) instruments and the LION (Low Energy Ion and
Electron Instrument) sensor on the COSTEP (Comprehensive Suprathermal and
Energetic Particle Analyzer) instrument were all successfully switched on
yesterday and today. Yesterdays first-light images from the MDI instrument
showed a notable blotchy pattern that was assumed to be related to
non-thermal equilibrium in the MDI filter. Today, the filter reached
equilibrium and new imagery from MDI show practically no trace of the
blotchiness observed yesterday.

     One of the most significant (and worrisome) events occurred today: the
activation and recommissioning of the EIT (Extreme ultraviolet Imaging
Telescope) instrument. EIT was successfully turned on today (13 October) and
four images from each of the four extreme ultraviolet wavelengths were
successfully obtained. An examination of the images revealed excellent image
quality. It appears the CCD detector has actually benefited from the 3 month
bakeout that occurred when the spacecraft lost control. Current
FULL-RESOLUTION imagery is now being released in honour of the spacecrafts
second year of operation. The imagery may be obtained at:
http://umbra.nascom.nasa.gov/eit/eit_full_res.html
Please note that these images are not true with respect to the Sun's north
and south pole. The orientation of the spacecraft has not yet been trimmed so
that these sensors see the Sun's north pole at the top of the image (as was
the case prior to the spacecraft problem). In these images, the northern
solar pole is more nearly toward the lower-right side of the image frames
while the eastern limb is toward the upper-right and the western limb is
toward the lower-left of the image frame. This oddity, which affects all of
the imaging instruments, will be corrected when the spacecraft is rolled into
an appropriate orientation.

     The LASCO (White Light And Spectrometric Coronagraph) instrument was
also successfully activated today. All heaters, calibration lamps and
shutters of each of the C1, C2 and C3 cameras have been tested successfully.
Dark images from each of the cameras at 0 seconds and 60 seconds have been
obtained. Activities currently in-progress include checking out the C2 and C3
filter wheels, mechanisms and the cameras themselves. This will be followed
by checking the C1 mechanisms and camera, and then by a complete calibration
sequence of the LASCO telescopes. Not until these steps have been completed
will the telescope doors be opened and the solar corona be imaged.

     The process of recommissioning the UVCS (UltraViolet Coronagraph
Spectrometer) instrument is also continuing without problems today.

     It has been a GREAT day for SOHO!



**  End of Bulletin  **

-- END included message